The homework my teacher assighned asked us to turn an equation in slope-intercept form into one in standard form. I have no problems with this, but it also asked us to find the x-intercept from the equation. Could you please explain to me how to find the x-intercept from a standard equation, please?
Answer by checkley75(3666)

in the standard line formula Y=mX+b you find the x intercept by setting the y value to 0 & solve for x thus:
Y=mX+b
mX=Y-b
X=y/m-b/m
EXAMPLE
Y=3X+5
0=3X+5
3X=-5
X=-5/3 WHIC IS THE X INTERCEPT (-5/3,0)